OpenStudy (anonymous):

a*x^(a-1)*y^b+z=0 b*x^(a)*y^(b-1)+z=0 x+y+1=0 solve for x and y

OpenStudy (anonymous):

Maybe you should write out the whole equation and not in this form... It's too messy...

OpenStudy (anonymous):

\[a x^{a-1}y ^{b}+\lambda=0\] \[bx ^{a}y ^{b-1}+\lambda=0\] x+y-1=0

OpenStudy (anonymous):

\[ \left\{ \begin{array}{l} ax^{a-1}y^b+z=0\\ bx^ay^{b-1}+z=0\\ x+y+1=0 \end{array}\right. \]

OpenStudy (anonymous):

Now is it x+y+1 or x+y-1 ?

OpenStudy (anonymous):

\[ \left\{ \begin{array}{l} ax^{a-1}y^b+z=0\\ bx^ay^{b-1}+z=0\\ x+y-1=0 \end{array}\right. \] Plug the second equation into the first and you get \[ x^{a-1}y^{b-1}(ay-bx) = 0 \] If \[a>1\] then x=0 and y=1 is a solution. If \[b>1\] then x=1 and x=0 is a solution. If \[a+b\not=0\] then you may infer by plugging the above equation into the third one that \[ x=\frac{a}{a+b}, y=\frac{b}{a+b} \].